# Retrieval-Augmented Generation for AI-Generated Content: A Survey

### v3

@@ IV-A2 Fact Verification
Fact Verification involves assessing the veracity of information, a critical function in disciplines such as Natural Language Processing (NLP), Information Retrieval, and Data Mining. In today’s digital era, characterized by an exponential increase in data, particularly across social media and online news platforms, there is a rapid proliferation of unchecked information. Fact verification plays an essential role in countering the spread of fake news, deceptive content, and rumors, thereby preserving the integrity of the information landscape and ensuring the public’s access to accurate knowledge. Consequently, automated fact verification systems are of immense importance, with broad applications and significant practical value. CONCRETE [230] leverages cross-lingual retrieval mechanisms to tap into a wealth of multilingual evidence, effectively bridging the gap in resources for languages that are underrepresented in fact-checking datasets. Hagström et al. [231] proved on LLaMA [4] and Atlas [30] that search augmentation is more beneficial for solving inconsistency problems than increasing model size. Atlas [30] shows that using RAG to support LLMs in knowledge-intensive tasks markedly improves their few-shot learning performance.
@@ IV-A3 Commonsense Reasoning
Commonsense Reasoning entails the capability of machines to infer or make decisions on problems or tasks in a human-like manner, drawing upon their acquired external knowledge and its application. However, the vast scope of common sense knowledge and the intricacies of reasoning processes make Commonsense Reasoning a perennially challenging and prominent area of research within the field of NLP. KG-BART [232] expands the conceptual landscape by incorporating intricate interrelations among diverse concepts within a knowledge graph. It employs graph attention mechanisms to aid LLMs in crafting more nuanced and logically coherent sentences. This approach not only improves the models’ generalization capabilities but also significantly bolsters their Commonsense Reasoning proficiency. Wan et al. [233] constructed the CONFLICTINGQA dataset, comprising contentious questions and conflicting answer documents, to examine which textual features significantly influence LMs’ ability to independently navigate controversial issues. The findings reveal that LMs often neglect the stylistic aspects of text that are typically valued by people.
@@ IV-A4 Human-Machine Conversation
Human-Machine Conversation encompasses the ability of machines to comprehend natural language and adeptly employ this skill to engage with humans seamlessly. This capability represents a significant challenge within the realms of Artificial Intelligence and Natural Language Processing and offers a broad spectrum of practical applications. As such, Human-Machine Conversation continues to be a focal point of research for many scholars. ConceptFlow [234] leverages a commonsense knowledge graph to structure conversations, directing the flow of dialogue based on attention scores, and propelling the conversation forward. This method achieves commendable results even with a substantial reduction in model parameters. Cai et al. [235] reimagined the text generation task as a cloze test by retrieving and distilling the essence of past conversational history, leading to notable outcomes. Komeili et al. [236] augmented dialogue generation quality by harnessing advanced search engine technologies to source pertinent content from the internet. BlenderBot3 [237] broadens its search horizon, not only mining relevant internet content but also local dialogue history, and employs entity extraction among other techniques to refine the quality of the resulting dialogue. Kim et al. [238], PARC [239], and CREA-ICL [240] improve the caliber of non-English conversations by incorporating cross-lingual knowledge, effectively addressing the scarcity of non-English datasets and enhancing the quality of the generated dialogue. CEG [241] addresses hallucination issues through a post-processing mechanism, verifying LLM-generated answers through retrieval. If the answer is correct, the retrieved document is added to the original answer as a reference; if the answer lacks reliable references, it guides the LLM to respond to the question anew.
@@ IV-A5 Neural Machine Translation
Neural Machine Translation (NMT) is the automated process of translating text from a source language to a target language [161, 242, 243]. It is a pivotal task in the domain of NLP and represents a significant objective in the pursuit of AI, boasting considerable scientific and practical significance. Cai et al. [242] proposed an innovative approach that utilizes monolingual corpora alongside multilingual learning techniques, challenging the traditional dependency on bilingual corpora in Neural Machine Translation. This approach ensures that the retrieval system provides ample information while simultaneously optimizing both the retrieval mechanism and the translation model, culminating in impressive performance. KNN-MT [243] executes translation tasks at the token level by computing vector space distances. TRIME [161] effectively minimizes the discrepancy between training and inference phases by jointly training the retrieval system and the generation model, thereby enhancing the precision of translations.
@@ IV-A6 Event Extraction
Event Extraction is a specialized task within Natural Language Processing (NLP) that focuses on pinpointing and extracting instances of particular event types from unstructured textual data. An event is generally characterized by a central action or predicate and the related entities, which can include participants, temporal indicators, locations, and other relevant attributes. The objective of event extraction is to convert the nuanced details embedded within text into a structured format, thereby facilitating advanced analysis, efficient information retrieval, and practical downstream applications. R-GQA [244] employs a retrieval-based approach to enhance the context of a given issue by identifying and utilizing the most closely aligned Question-Answer pair from a repository, thereby enriching the information available for processing the current query.
@@ IV-A7 Summarization
In the realm of NLP, Summarization is a task aimed at distilling the essential information from lengthy texts and producing a concise, coherent summary that encapsulates the primary themes. Summarization enables users to quickly grasp the essence of a text, thereby conserving time that would otherwise be spent on reading extensive material. There are two main approaches to Summarization: Extractive and Abstractive. Extractive Summarization involves the automatic selection and compilation of key phrases directly from the source text. A key phrase succinctly captures the main themes, content, or perspectives of the text and is typically composed of one or several words. The generation of key phrases is instrumental for understanding, categorizing, retrieving, and organizing textual information. It is extensively applied in fields such as search engine optimization, academic research, text summarization, and more. This technique refrains from creating new sentences, instead repurposing segments from the original text. Abstractive Summarization, on the other hand, entails comprehending the original text’s meaning and reformulating it into new sentences [245, 153, 246, 247]. This approach can convey the source’s intent more fluidly but poses greater challenges in terms of implementation due to its complexity. RAMKG [245] effectively leverages a comprehensive English corpus to bolster the performance of Keyphrase Generation in non-English contexts. It does so by enhancing the alignment of keywords extracted from texts in different languages that share similar subject matter. Unlimiformer [153] addresses the issue of input length constraints in transformer-based models by retrieving and utilizing the top-k most relevant hidden states, thereby extending the model’s capacity to handle longer inputs. RPRR [246] employs a Retrieve-Plan-Retrieve-Read approach to overcome the limited context window constraints faced by LLMs, utilizing retrieved information to generate high-quality Wikipedia documents for emerging events. RIGHT [247] chooses to use different types of retrievers in different datasets to enhance the generator, which can effectively improve the quality of automatically generated labels in simple deployment.
@@ IV-B1 Code Generation
The goal of code generation is to transform natural language (NL) descriptions into code implementation, which can be seen a process of text-to-code. Therefore, LSTM and transformer models are widely used for generator. Whether to use code-specific retrieval or text-based retrieval depends on the contents to be searched.
Retrieval-based prompt engineering is one of the most prevalent scenarios of RAG in code generation. In-context learning includes training samples in prompts as the input for sequence-to-sequence generative models. Retrieval techniques are adopted to find similar training samples to the test input, so that the prompt can be more informative and related. REDCODER [40] retrieves similar NL descriptions using dense retriever CodeBert [25], then concatenates the NL texts, their paired codes, for downstream generator PLBART [41]. CodeT5Mix [248] adopts the paradigm of RECODER, proposing a model that functions dually as both the retriever and generator. APICoder [129] first train a Bert-based deep retriever to align the embeddings of NL descriptions and API documentation; then, it retrieves relevant API information to build prompt for the generator CODEGEN-MONO [249]. The following work [250] uses the same pipline, renaming the retriever module as APIFinder. COCOGEN [251] aims at commonsense reasoning, which generates code-based reasoning-graphs with Codex [2] given NL inputs; it adds an evaluation setting of dynamic prompt selection, which actually retrieves relevant examples for prompts. In DocPrompting [42] given an NL intent, the retriver retrieves relevant documentations, then the generator generates codes based on the NL and retrieved documents. It evaluates both sparse retrievers and dense retrievers, and also tries different generators in experiments. CodeT5+ [252] adopts the CodeT5 [106] model for both the retriever and generator, leveraging only the encoder part in the retrieval process. AceCoder [188] fixes the retriever to BM25 [19], and tests several LLM generators for code generation. A3CodGen [253] extracts local information, retrieves relevant global functions using embeddings, and incorporates third-party library information to construct the prompt for LLM-based code generation. SKCODER [254] employs BM25 to retrieve relevant code snippets, which are further processed to produce sketch template. The template and the original description are concatenated for final generation. CodeGen4Libs [255] uses RAG for both import statements and codes, employing BM25 as retriever and finetuned CodeT5 as generator. CODEAGENT [256] design agents to search on web, retrieve relevant documentation, generate programs, and test correctness. RRGCode [257] retrieves relevant code snippets using both sparse and dense retrieval, employs a cross-encoder to re-rank the retrieval results, then generates code using the concatenation of the query and the retrieved codes. A recent study [258] shows that retrieval-augmented framework for code suggestions, including code generation and code completion, can improve the performance by a large margin. ARKS [259] steps further upon prompt-based RAG, incorporating iterative RAG to re-formulate queries and update knowledge soup (containing documentation, execution feedback, generated code, etc.) for dense retrieval, improving final generation accuracy.
Retrieval results can be applied during the generation process as well. RECODE [120] retrieves NL descriptions and paired codes using edit distance, then extracts n-gram action subtrees from codes’ ASTs. During LSTM-based generation, the patterns of processed subtrees are leveraged to increase the corresponding word probability at each decoding step. kNN-TRANX [163] uses seq2tree model BertranX [260] to convert NL to code AST. It constructs a datastore for each code AST prefix and NL pair; i.e., for each NL-code pair, the context representation of the i-th context is obtained by encoding NL and the i-th prefix of code’s AST through the seq2tree model. At each decoding step of the generation, the hidden representations are searched within the datastore to form a new probability, which is later combined with the seq2tree model’s output through a confidence network.
ToolCoder [261] performs normal code generation, and conducts online search or offline retrieval when encountering special <API><API> token. This paradigm makes the model learn to leverage API tools.
@@ IV-B2 Code Summarization
The goal of code summarization is to transform code into natural language (NL) descriptions, which is a process of code-to-text. Same to code generation, many sequence-to-sequence models are applied as generator.
In many research works, the retrieval results are processed by additional encoders.
@@
Re2Com [141] and EditSum [138] both retrieve similar code using sparse retrieval BM25 and generate summary using LSTM. As shown in Fig. 8, they separately encode the input, the retrieved code, and the corresponding summary, then combine the middle representations or probabilities in the decoder. HGNN [262] instead uses code edit distance for retrieval, and substitutes the encoders for codes with hybrid GNN on their Code Property Graphs (CPG) [263]. RACE [142] aims at generating commit message for code difference. It leverages dense retrieval for similar examples and transformer model for generation. It also uses three encoders for the input, the retrieved code difference, and corresponding commit message, then combines the results before feeding into the decoder. BASHEXPLAINER [139] shares the similar idea. Its dense retrieval module is based on CodeBert [25], and for generation, the output representations of the input and the similar code from CodeBert are directly fused for transformer-based decoder. READSUM [264] retrieves similar code using Levenshtein distance, applies code encoder and summary encoder to retrieved pairs, and generates summary using a decoder where a fusion network combines the information.
RAG for in-context learning, which retrieves similar examples and build prompt for generation, also works in code summary. REDCODER [40] works for both code generation and summary, and it replaces the retriever with GraphCodeBert [105] for code summary task. ASAP [203] retrieves similar code with BM25, and generates summary with GPT models. Similar to the paradigm described above, research on pseudocode generation [265] employs the retrieved code as input for generation and subsequently replaces the results with the original input. SCCLLM [266] retrieves relevant code snippets by semantic, syntactic, and lexical-based retrieval, then forms in-context prompts for smart contract comment generation via LLM. UniLog [267] retrieves similar code snippets paired with their log statements to conduct in-context learning for log statement generation.
Logit-based RAG also prevails in code summarization. Rencos [121] utilizes two different methods to retrieve similar code snippets, syntactic similarity between abstract syntax trees (AST) and cosine similarity between dense representations. For generator, it adopts attention-based LSTM. There are three encoder-decoder models for the code input and two retrieval results respectively, and the probabilities are combined for final generation. CoRec [268] generates commit message given code diff and retrieved similar code diff. Multiple LSTM generators handle the input code diff and the retrieved code diff, then adds the probabilities for final generation. kNN-Transformer [269] obtains context vector by feeding code into an encoder-decoder generator, then gets three parts of logits: the first is from searching the vector, the second is from normal Transformer, and the third is the copy mechanism that reserve rare tokens from the input. Tram [270] involves retrieval in both token-level and sentence-level. It encodes source code and corresponding AST into hidden states representations; for token-level, it retrieves similar representations in the database to form next-token prediction logits; for sentence-level, it uses similar code for autoregressive generation logits; it also add the original autoregressive generation logits. CMR-Sum [271] uses encoder-decoder model to generate summaries. It conducts cross attention between representations of retrieved summary and generated summary, and add the logits to the original transformer logits for final distribution.
@@ IV-B3 Code Completion
Code completion can be thought of as the coding equivalent of the “next sentence prediction” task. Early attempts on function completion [272] employs DPR to retrieve relevant template functions using function docstring, then concatenate the information as the input to the code generator BART. ReACC [132] retrieves similar codes to build prompts for generation. For retrieval, it uses hybrid retriever, which combines sparse and dense retrieval; for generation, it uses CodeGPT-adapted [273]. RepoCoder [218] steps further to perform iterative retrieval and generation to bridge the gap between the retrieval context and the intended completion target. In each iteration, for retrieval, the code query is augmented with previously generated code; for generation, the prompt is formed by combining the newest retrieved codes with the code query. Other than combining retrieval results in prompts, the retrieval-and-edit framework [140] first retrieves similar training examples using dense retrieval, then encodes the input and the retrieved result separately, finally combine them through attention for later LSTM decoder. CoCoMic [274] builds a project context graph for the whole code project, and retrieves top-k neighbors of the input source code. It generates representations of both source code and retrieved contexts, then jointly processes the embeddings to complete the code. RepoFusion [275] follows the idea of Fusion-in-Decoder; it employs multiple encoder to input the concatenation of the retrieved repo contexts and the unfinished code, them fuses them and generates the results through a decoder. KNM-LM [276] uses the same model for retrieval encoding and generation, then combines the logits using bayes inference. EDITAS [277] aims at assertion generation. It retrieves similar queries and their assertions, encodes the edit sequence (from the retrieved query to the original query) and the retrieved assertion, then fuses the information for decoder generation. De-Hallucinator [278] first generates code snippets without retrieval, then retrieves relevant API references given generated contents. In the second pass, retrieved API references are combined in prompt for better generation. REPOFUSE [279] uses both rationale context and retrieved similar codes to construct prompt. To fit in the input length limit, it reserves the contexts that are most relevant to the query.
@@ IV-B4 Automatic Program Repair
Buggy code can take a lot of effort to fix. Automatic program repair leverages generative models to output the correct version. Query-based RAG techinque is widely used in automatic program repair [209, 130, 131, 182, 208]. Among them, RING [209] retrieves similar error messages based on both sparse and dense vectors, then builds prompt for the generator Codex [2]. CEDAR [130] applies for both assertion generation and program repairs tasks; it uses sparse and dense retrieval to search for similar codes, then forms prompt for Codex to generate results. InferFix [131] crafts a prompt carrying the bug type, location, relevant syntax hierarchies, and similar fixes through dense retrieval. Then it also uses Codex for generation. RAP-Gen [182] also retrieves similar buggy codes and corresponding fixes through hybrid retriever (including both sparse and dense retriever). It finetunes CodeT5 [106] with this RAG paradigm. SARGAM [208] searches similar buggy code using dense retrieval, generates patches using augmented input, then applies another model to modify the final result. These research works also involve error localization, which is not our focus. RTLFixer [280] leverages ReAct and RAG to implement an agent fixing errors in Verilog codes. It iteratively retrieves relevant errors and corresponding solutions, and combines reasoning and action planning to form prompts for LLMs.
@@ IV-B5 Text-to-SQL and Code-based Semantic Parsing
Semantic parsing is the task of translating natural language utterances to unambiguous structured meaning representations, where code language is often leveraged to augment this process. SQL is not only a programming language but can also be considered as a structured representation, so we place text-to-SQL (a special case of code generation) in this subsection. Related research works all apply query-based RAG. XRICL [189] focuses on the problem of translating non-English utterances into SQL queries. It searches and reranks English utterance using non-English ones by dense retrieval, then builds prompt for Codex to generate SQL queries. SYNCHROMESH [122] proposes constrained semantic decoding to enforce rich syntactic and semantic constraints during generation of SQL or other languages. It uses the similarity between abstract syntax trees (AST) to finetune the dense retriever S-Bert. During inference, similar NL and SQL are retrived to form the prompt of GPT-3. CodeICL [281] uses Python for semantic parsing, and augments prompts with a structured domain description for GPT generation. In few-shot learning setting, it leverages BM25 sparse retrieval to find similar training examples. RESDSQL [282] ranks schemas using cross-encoder, then includes ranked schemas into prompts to generate SQL skeleton and SQL query. ReFSQL [283] retrieves relevant questions and corresponding SQL to augment text-to-SQL generation. It involves structure-enhanced retriever with schema linking, and Mahalanobis contrastive learning to improve the retrieval performance. ODIS [284] retrieves in-domain and out-of-domain demonstrations using BM25, then includes them for in-context learning to generate SQL. Another work [285] retrieves both similar and diverse demonstrations, and then builds prompt for in-context learning to generate SQL. MURRE [286] conducts multi-hop retrieve-rewrite, where relevant tables are retrieved through dense retrieval and then re-writed to generate new tabularized question. A rank module at last integrate the retrieval results and select the top tables for the text-to-SQL module. CodeS [287] retrieves relevant information from table databases in a coarse-to-fine manner, then includes retrieved values to build prompts for both finetuning and inference.

### v4

@@ IV-A2 Fact Verification
Fact Verification typically refers to determining whether a given natural language text and a related claim or assertion match the facts in the text. CONCRETE [187] leverages cross-lingual retrieval mechanisms to tap into a wealth of multilingual evidence, effectively bridging the gap in resources for languages that are underrepresented in fact-checking datasets. Hagström et al. [188] proved on LLaMA [4] and Atlas [30] that search augmentation is more beneficial for solving inconsistency problems than increasing model size. Atlas [30] shows that using RAG to support LLMs in knowledge-intensive tasks markedly improves their few-shot learning performance.
@@ IV-A3 Commonsense Reasoning
Commonsense Reasoning entails the capability of machines to infer or make decisions on problems or tasks in a human-like manner, drawing upon their acquired external knowledge and its application. KG-BART [189] expands the conceptual landscape by incorporating intricate interrelations among diverse concepts within a knowledge graph. It employs graph attention mechanisms to aid LLMs in crafting more nuanced and logically coherent sentences. Wan et al. [190] constructed the CONFLICTINGQA dataset with contentious questions and conflicting answers to study how textual features affect LMs’ handling of controversial issues.
@@ IV-A4 Human-Machine Conversation
Human-Machine Conversation encompasses the ability of machines to comprehend natural language and adeptly employ this skill to engage with humans seamlessly. ConceptFlow [191] leverages a commonsense knowledge graph to structure conversations, directing the flow of dialogue based on attention scores, and propelling the conversation forward. Cai et al. [192] reimagined the text generation task as a cloze test by retrieving and distilling the essence of past conversational history, leading to notable outcomes. Komeili et al. [193] augmented dialogue generation quality by harnessing advanced search engine technologies to source pertinent content from the internet. BlenderBot3 [194] broadens its search horizon, not only mining relevant internet content but also local dialogue history, and employs entity extraction among other techniques to refine the quality of the resulting dialogue. Kim et al. [195], PARC [196], and CREA-ICL [197] improve the caliber of non-English conversations by incorporating cross-lingual knowledge, effectively addressing the scarcity of non-English datasets and enhancing the quality of the generated dialogue. CEG [198] addresses hallucination issues through a post-processing mechanism, verifying LLM-generated answers through retrieval.
@@ IV-A5 Neural Machine Translation
Neural Machine Translation (NMT) is the automated process of translating text from a source language to a target language [117, 199, 200]. It is a pivotal task in the domain of NLP and represents a significant objective in the pursuit of AI, boasting considerable scientific and practical significance. Cai et al. [199] proposed an innovative approach that utilizes monolingual corpora alongside multilingual learning techniques, challenging the traditional dependency on bilingual corpora in Neural Machine Translation. KNN-MT [200] executes translation tasks at the token level by computing vector space distances. TRIME [117] effectively minimizes the discrepancy between training and inference phases by jointly training the retrieval system and the generation model, thereby enhancing the precision of translations.
@@ IV-A6 Event Extraction
Event Extraction is a process in NLP that involves identifying and categorizing specific events within a text and associating them with relevant entities. These events are usually represented by verbs and the entities are the participants involved in the event. R-GQA [201] enhances the context of a given issue by identifying and utilizing the most closely aligned Question-Answer pair from a repository, thereby enriching the information available for processing the current query.
@@ IV-A7 Summarization
Summarization is a task aimed at distilling the essential information from lengthy texts and producing a concise, coherent summary that encapsulates the primary themes. There are two main approaches to Summarization: Extractive and Abstractive. Extractive Summarization involves the automatic selection and compilation of key phrases directly from the source text. This technique refrains from creating new sentences, instead repurposing segments from the original text. Abstractive Summarization, on the other hand, entails comprehending the original text’s meaning and reformulating it into new sentences [202, 95, 203, 204]. This approach can convey the source’s intent more fluidly but poses greater challenges in terms of implementation due to its complexity. RAMKG [202] effectively leverages a comprehensive English corpus to bolster the performance of Keyphrase Generation in non-English contexts. Unlimiformer [95] addresses the issue of input length constraints in transformer-based models by retrieving and utilizing the top-k most relevant hidden states, thereby extending the model’s capacity to handle longer inputs. RPRR [203] employs a Retrieve-Plan-Retrieve-Read approach to overcome the limited context window constraints faced by LLMs, utilizing retrieved information to generate high-quality Wikipedia documents for emerging events. RIGHT [204] chooses to use different types of retrievers in different datasets to enhance the generator.
@@ IV-B1 Code Generation
Code generation aims to convert Natural Language (NL) descriptions into code implementations.
Query-based RAG is a common method for code generation. It builds prompts for transformer-based generative models with retrieved information, including similar examples [40, 205, 206, 207, 144, 208], relevant API details [87, 209], documentations [42], imports [210], and global functions [211]. SKCODER [212] retrieves relevant code snippets to produce sketch template for final code generation. RRGCode [213] employs a cross-encoder to rank the retrieval results. CODEAGENT [214] designs agents for web search, documentation retrieval, program generation, and correctness testing. ARKS [215] incorporates iterative RAG to re-formulate queries and update retrieval sources.
Logit-based RAG is also applicable for code generation. RECODE [78] retrieves NL descriptions and paired codes using edit distance, then extracts n-gram action subtrees from ASTs. During LSTM-based generation, the processed subtrees are leveraged through logits at each decoding step. kNN-TRANX [119] uses a seq2tree model to convert NL to code AST. At each decoding step of the generation, the hidden states are searched within the AST prefix datastore to form new probabilities, which are later combined with the seq2tree model’s output through a confidence network.
ToolCoder [216] generates codes containing special tokens. When it encounters these tokens, ToolCoder performs online search or offline retrievals to fill in the blanks with API calls, which is a specialized form of speculative RAG.
@@ IV-B2 Code Summarization
Code summarization tasks in turn convert the code into NL descriptions.
Many research works process retrieval results using additional encoders and then combine them for subsequent decoder, which is similar to the Fusion-in-Decoder [35].
@@
Re2Com [100] and EditSum [97] retrieve similar codes using BM25 and generate summary using LSTM. They separately encode the input, the retrieved code, and the corresponding summary, then combine the hidden states or logits in the decoder. HGNN [217] instead uses code edit distance for retrieval, and substitutes the code encoder with hybrid GNN on their Code Property Graphs [218]. RACE [101] employs separate encoders for the input code difference, the retrieved code differences through dense retrieval, and corresponding commit message to generate the final commit messages. BASHEXPLAINER [98] applies dense retrieval, and fuses the embeddings for subsequent transformer-based decoder. READSUM [219] uses Levenshtein distance for retrieval, and employs a fusion network to combine the representations of retrieved codes and summaries.
Query-based RAG is prevalent for code summary generation. REDCODER [40], ASAP [160], and SCCLLM [220] all form prompts with retrieved contents for summarization. They employ dense retrieval, sparse retrieval, and hybrid retrieval (including semantic, syntactic, and lexical-based retrieval), respectively. The paradigm is also leveraged for pseudocode generation [221] and log statement generation [222].
Logit-based RAG also prevails in code summarization. Rencos [79] and CoRec [223] retrieve similar code snippets or code differences through AST or dense representations. They both adopt multiple LSTMs for the input and the retrieved results, and the probabilities are combined for final generation. kNN-Transformer [224] uses a transformer-based generator to obtain context vectors of input codes, then combines three parts of logits from vector search, the generator, and the copy mechanism for rare tokens in the input. Tram [225] also combines three sets of logits from the original generator, the generator for sentence-level retrieved results, and the search logits of the token-level vectors (which represent the source codes and their ASTs). CMR-Sum [226] incorporates the cross-attention probabilities between the retrieved summary and the generated summary, to the original generation logits.
@@ IV-B3 Code Completion
Code completion is akin to the code version of the “next sentence prediction” task.
Query-based RAG is the mainstream paradigm for code completion. Drain et al. [227] retrieved template functions for function completion. ReACC [90] uses both sparse and dense retrieval. RepoCoder [175] performs iterative RAG by augmenting the retrieval input with previously generated code. De-Hallucinator [228] retrieves API references using first-time generated contents, then conducts query-based RAG for improved code completion. REPOFUSE [229] includes rationale context and retrieved codes to form prompt, and ranks the contexts to fit in the length limit.
Many works leverage latent representation-based RAG. Retrieve-and-edit [99], RepoFusion [230], and EDITAS [231] employ multiple encoders for retrieved contents or edit sequences, then fuse the information for subsequent decoder. CoCoMic [232] retrieves codes on the project context graph of the whole code project. It jointly processes the representations of source codes and retrieved contexts in the generator.
KNM-LM [233] performs logit-based RAG, combining the logits of retrieval and generation using bayes inference.
@@ IV-B4 Automatic Program Repair
Query-based RAG is often used in automatic program repair to help generative models fix buggy codes. RING [166], CEDAR [88], and RAP-Gen [138] all use hybrid retrieval (including both sparse and dense retrieval) for similar error messages, buggy codes, or fixes to build prompts. InferFix [89] includes the bug type, the location, relevant syntax hierarchies, and similar fixes into the prompt. SARGAM [165] utilizes prompts with similar buggy codes to generate patches; then another model is employed to refine the final result. RTLFixer [234] leverages ReAct [128] to implement an agent fixing errors in Verilog codes. It iteratively retrieves errors and paired solutions, and combines reasoning and action planning into prompts for LLMs.
@@ IV-B5 Text-to-SQL and Code-based Semantic Parsing
Semantic parsing converts NL into clear, structured representations, like SQL or other domain-specific languages, often with the assistance of codes. All related works that employ RAG specifically utilize its query-based variant. XRICL [145] searches and reranks English utterance using non-English ones, then builds prompt to generate SQL queries. SYNCHROMESH [80] retrieves similar NL and SQL to build prompts, then conducts constrained semantic decoding to enforce rich syntactic and semantic constraints during SQL generation. CodeICL [235] uses Python for semantic parsing, leveraging BM25 to incorporate similar training examples into prompts. RESDSQL [236] includes ranked schemas into prompts to generate SQL skeleton and SQL query. ReFSQL [237] uses a structure-enhanced retriever with schema linking and Mahalanobis contrastive learning, which helps to make better text-to-SQL generation. To build prompts for SQL generation, ODIS [238] retrieves both in-domain and out-of-domain demonstrations, while Nan et al. [239] retrieved both similar and diverse demonstrations. MURRE [240] conducts multi-hop retrieve-rewrite on tables to generate tabularized question, then ranks the results for prompt construction. CodeS [241] retrieves relevant information from table databases in a coarse-to-fine manner to generate SQL.